a variable force f acts on a body which is free to move the displacement of the body is proportional to t3 then power delivered is proprtional to

Grade:11

1 Answers

mukul chandrakar
6 years ago
Force= m.aForce proportional to adifferentiate a(t^3) we get velocity =3t^2 power = F.v by proptionainty (a.v) =t^5